# Adversarial Training for Failure-Sensitive User Simulation in Mental Health Dialogue Optimization

### v2

@@ 4.2 Model Configurations
We train simulators by fine-tuning Llama-3.3-70B-Instruct using LoRA (hu2022lora) on datasets of varying sizes (10K, 20K, 40K, 80K, 160K sessions) to explore data scale effects. We perform adversarial training on the final checkpoint (160K), producing UserSim-160K-it1, UserSim-160K-it2, and UserSim-160K-it3 through successive refinement iterations. For comparison, we evaluate three zero-shot base models: GPT-4o (openai2024gpt4o), Kimi-K2-Instruct (kimiteam2025k2), and Llama-3.3-70B-Instruct. Training hyperparameters are provided in Appendix C.2.
@@
For each configuration, we generate 2,000 simulated sessions: 1,000 unique session contexts drawn from real user profiles and conversation histories × 2 simulations per context, with early stopping when conversations naturally conclude. Our context construction for user simulation follows the hierarchical structure described in Appendix A, including user profiles, multi-session history, and current session grounding. To isolate simulator quality effects, we use a fixed set of 14 chatbot models to simulate conversations.
@@ 4.3.1 Linguistic Features
Type-Token Ratio (TTR): The ratio of unique words (types) to total words (tokens) in a session, measuring lexical diversity. TTR ranges from 0 to 1, with higher values indicating greater vocabulary variety.
@@ 4.3.2 Behavioral Metrics
User Action Distribution: Conversations are annotated with mental health-relevant user actions using GPT-5-nano to classify each individual user message according to our action taxonomy (see Appendix D for example categories). Since individual messages can have multiple actions, the action distributions are not normalized (probabilities do not sum to 1).
Issue Rate Analysis: A critical requirement for realistic user simulation is the ability to expose chatbot failure modes. We evaluate simulators by measuring how well they replicate the rate and distribution of problematic conversational features (“issues”) that chatbots exhibit with real users. Issues are identified using GPT-5 with low reasoning effort to analyze entire sessions. The analyzer identifies issues from a detailed taxonomy (35+ categories including invalidating responses, unclear wording, hallucinations, etc.) and highlights the specific messages where each issue occurred, enabling proper credit assignment (see Appendix E for example categories). We compare issue rates and category distributions between simulated and real conversations to validate that our simulator accurately reflects the chatbot’s real-world performance and can reliably identify system weaknesses before deployment. Since we examine proportions of issue categories (which form a valid probability distribution that sums to 1), we use KL divergence to measure distributional alignment:
@@
Similar to the discriminator, the issue analyzer examines whole sessions rather than isolated messages. This design is essential because messages are only problematic when considered in context with the user’s situation and their subsequent replies. The session-level approach makes realistic user simulation critical for accurate evaluation—the simulator must produce contextually appropriate responses that reveal when the chatbot’s messages fail to address the user’s actual needs. This enables numerous evaluation dimensions that would be impossible with isolated message analysis: detecting circular conversations, identifying premature topic changes, recognizing when the chatbot ignores user requests, and assessing whether mental health progress occurs naturally over the conversation arc.
@@ 4.3.3 Discriminator Performance
We measure discriminator accuracy using fine-tuned discriminators trained on balanced datasets of real and simulated sessions from their corresponding simulator versions. Each discriminator is trained under identical conditions—1,600 training sessions and 400 validation sessions, with the same learning rate and number of epochs—ensuring that accuracy differences reflect simulator quality rather than training differences. Lower discriminator accuracy indicates a more realistic simulator, with the target being 50% (random guessing).
@@ 5 Results
Fine-tuned simulators dramatically outperform zero-shot models across all metrics. t-SNE visualization (Appendix B.1) shows that embeddings of conversations simulated by base models form distinct clusters separate from their real counterparts. The fine-tuned simulators’ outputs, on the other hand, form tight clusters defined by shared context rather than by the generation process, demonstrating effective grounding. Zero-shot models produce unrealistically compliant users, as evidenced by significantly lower issue rates (Figure 2) and severe overrepresentation of positive mental health indicators (Appendix B.3).
Table 1: Linguistic diversity features comparing real users with simulators. Entropy measures vocabulary diversity; TTR (type-token ratio) measures lexical diversity; Diversity measures bigram diversity. Values closer to real users indicate better linguistic realism. Best fine-tuned values (closest to real users) are in bold.
@@
As shown in Table 1, which compares linguistic features across simulators, some base models achieve diversity metrics closer to real users than fine-tuned models (e.g., Kimi-K2-Instruct’s bigram diversity of 0.851 vs. real users’ 0.861) as supervised fine-tuning exhibits non-monotonic effects on lexical diversity. Bigram and TTR initially improve from 10K-40K samples as the model learns to use conversational context to improve response diversity. However, mode collapse emerges from 40K-160K samples as continued training learns common patterns at the expense of rare variations. For behavioral alignment, fine-tuned simulators dramatically outperform zero-shot base models, achieving total BCE losses of .159-.160 compared to .187-.207 for zero-shot models (see Appendix B.3). To evaluate how well simulators surface chatbot issues, we measure issue rates (percentage of chatbot messages containing issues) using GPT-5 to analyze complete sessions. Figure 2 shows that fine-tuned simulators produce rates of 3.07-3.31%, closely matching real conversations (3.02%), while zero-shot models show substantially lower rates of 0.98-2.54%. UserSim-160K-it2 achieves the closest calibration to the real rate (3.07%). Detailed issue category discrepancies are provided in Appendix B.2.
@@
5.2 Adversarial Training Improvements
Three iterations of adversarial training produced progressive improvements:
@@ 5.2 Adversarial Training Improvements
Linguistic Diversity: Adversarial training successfully addresses mode collapse by recovering lexical diversity while maintaining behavioral alignment. UserSim-160K-it3 progressively approaches real user levels across all diversity metrics, demonstrating that adversarial training addresses a fundamental limitation of supervised fine-tuning. However, oscillatory behavior across iterations indicates the need for careful hyperparameter tuning. We employ conservative dataset filtering (retaining only high-confidence preference pairs) to mitigate over-correction and maintain distribution stability.
Distributional Alignment: Adversarial training produces modest improvements in action distribution alignment (see Appendix B.3), with progressive refinement across iterations. The limited magnitude of improvements suggests the base fine-tuned model already achieved strong behavioral alignment.
Discriminator Fooling: Adversarial training dramatically reduced discriminator accuracy across conversation turns (Table 2). All discriminators were trained under identical conditions with on-policy optimization—each trained exclusively on sessions from its corresponding simulator version. The dramatic performance degradation confirms that adversarial iterations remove systematic artifacts and produce strong distributional alignment throughout full sessions.
Predictive Validity: To evaluate how well simulators enable reliable offline evaluation, we measure correlation between simulated and real issue rates across diverse chatbot models, and KL divergence between simulated and real issue category distributions. Table 3 shows that adversarial training produces substantial improvements in both metrics. UserSim-160K-it2 achieves the strongest correlation ($r=0.818$), while UserSim-160K-it3 achieves the lowest distributional divergence (KL = 0.148). This provides compelling evidence that adversarially trained simulators are significantly more realistic and effective at identifying chatbot problems in offline testing compared to their base fine-tuned counterparts. Notably, aggregating predictions across all simulator models achieves a strong correlation ($r=0.801$) and low distributional divergence (KL = 0.152), suggesting that ensembling multiple simulators could provide even more robust evaluation.
@@
A/B Test Replication: Beyond aggregate correlation, we validate predictive utility by replicating a recent model A/B test. All three models exhibited consistent directional trends between simulation and real data, and no model that failed in simulation improved in the wild (Appendix B.5). Combined with strong correlation and distributional alignment, these results validate the simulator’s utility for reliable offline evaluation.
@@ 6 Discussion
Firstly, zero-shot models exhibit severe distributional biases, producing unrealistically compliant users who systematically underrepresent negative and resistant behaviors. This confirms that generic language models, despite their capabilities, cannot reliably replicate domain-specific user behavior without exposure to real conversational patterns.
While supervised fine-tuning achieves strong behavioral alignment, it suppresses lexical diversity through mode collapse. Adversarial refinement recovers this diversity while maintaining behavioral fidelity—successfully producing simulators that fool increasingly sophisticated discriminators without sacrificing predictive utility. This demonstrates that realism requires both distributional accuracy and surface-level variation.
Lastly, unlike traditional metrics that measure task success or correlation, discriminator accuracy directly quantifies distributional realism by testing whether conversations are distinguishable from real interactions. The dramatic accuracy degradation across adversarial iterations provides evidence of improved simulation fidelity that complements task-level validation.
@@ 7 Conclusion
We have presented an adversarial training framework for creating realistic user simulators that can reliably expose failure modes in multi-turn task-oriented dialogue systems. Applied to mental health support—a complex TOD domain where users pursue mental health goals through extended goal-directed conversations—our approach demonstrates that domain-specific fine-tuning combined with adversarial refinement produces simulators that closely replicate real user behavior across linguistic, semantic, and behavioral dimensions.
Our key contributions are: (1) demonstration that fine-tuned models vastly outperform zero-shot base models for behavioral fidelity, (2) an adversarial training method that recovers lexical diversity suppressed by supervised learning while improving distributional alignment and predictive validity, and (3) comprehensive validation that adversarially-trained simulators reliably predict real model performance, achieving strong correlation with real-world results and successfully replicating A/B test trends.
These results provide evidence that adversarial training is a promising methodology for creating realistic user simulators in mental health support TOD domains. By enabling rapid iteration and systematic exploration of system failure modes before deployment, such simulators can reduce development time, protect users from suboptimal experiences, and potentially provide safe environments for reinforcement learning based optimization of conversation quality.
@@
Mode Collapse: While adversarial training substantially improves lexical diversity, fine-tuned simulators still show lower phrase-level variety than real users. Techniques like controlled generation or mixture-of-experts architectures might further improve diversity preservation.
Domain Transfer: Our evaluation focuses on mental health support—a complex, open-ended TOD domain. Validating the methodology in traditional TOD domains (restaurant booking, customer service) and other complex domains (tutoring, technical support) would establish broader applicability and identify domain-specific challenges.
RL Training Integration: While our simulators enable valid offline evaluation, integrating them directly into RL training loops for TOD policy optimization remains future work. Recent work on RL-based tutoring (nam2025efficient) demonstrates that simulator quality is critical for reliable RL-based dialogue optimization, underscoring the importance of adversarial training for creating simulators that support both evaluation and training.
@@
All data used in this research were collected from real user conversations with our mental health chatbot with explicit user consent. Personally identifiable information was removed through automated script-based detection and removal of names, phone numbers, email addresses, and other identifiers in accordance with data protection policies and ethical guidelines for human subjects research. However, due to the sensitive nature of mental health data, we are unable to share the datasets used in this research, despite our desire to promote reproducibility and open science.
Realistic user simulators could potentially be misused to train systems optimized for engagement over user wellbeing, or to manipulate vulnerable populations. Our simulators are designed for system evaluation and development, enabling rapid iteration without exposing users to suboptimal experiences. We note that simulators trained on real conversations may perpetuate biases present in the original dataset, potentially disadvantaging certain demographic groups. We emphasize that simulators should be deployed with appropriate oversight and ethical review. Additionally, while simulators can reduce the need for human testing in many scenarios, they should complement rather than entirely replace human feedback for validating novel interventions or edge cases.
@@ B.3 User Action Distributions
For positive indicators (right), zero-shot models exhibit the opposite bias—substantially overrepresenting “Change Talk” ($\sim$40% vs. $\sim$15% in real data) and “Agrees with Chatbot.” This creates an artificially positive mental health support environment that fails to challenge the chatbot or surface issues. Fine-tuned simulators (UserSim variants) achieve close alignment with real sessions across both indicator types, with adversarial training iterations showing incremental improvements in matching real distributions.
@@ B.4 Correlation with Production
The adversarial training process substantially improves correlation with real issue rates, as shown in Table 3. UserSim-160K-it2 achieves $r=0.818$ compared to UserSim-160K-base’s $r=0.528$, demonstrating the value of adversarial refinement. However, we note that several models achieve strong correlations even without adversarial training. This suggests that for rolled-out evaluation or RL training applications, ensembling multiple user simulator models could reduce noise and improve stability. Indeed, aggregating across UserSim-160K-base, it1, it2, and it3 achieves robust correlation ($r=0.801$) with low distributional divergence (KL = 0.152). The plot shown here highlights the best-performing model (UserSim-160K-it2) to demonstrate the upper bound of achievable predictive validity.
@@ C.2 Training Hyperparameters
Table 6 shows the hyperparameters used for supervised fine-tuning (SFT) of the base user simulator and Direct Preference Optimization (DPO) during adversarial training iterations.
@@ Appendix D User Action Categories
This appendix provides example user action categories from our taxonomy used to annotate and analyze user behavior in mental health support conversations. Each user message is classified using GPT-5-nano. Actions are classified as positive mental health indicators (Good), negative indicators (Bad), or user features that should change over time. The full taxonomy contains 20+ categories; we highlight key examples below.
@@ Appendix E Issue Categories
This appendix provides example issue categories from our taxonomy used to evaluate chatbot quality. Issues are identified by GPT-5 analyzing entire conversation sessions with low reasoning effort. For each detected issue, the analyzer highlights the specific affected messages to enable proper credit assignment. The full taxonomy contains 35+ categories organized into six main categories (User Distress, User Confusion, User Frustration, User Losing Trust, Chatbot Violates Policies, App Bugs); we highlight key examples below.

### v3

@@ 4.2 Model Configurations
We train simulators by fine-tuning Llama-3.3-70B-Instruct using LoRA (Hu et al., 2022) on datasets of varying sizes (10K, 20K, 40K, 80K, 160K sessions) to explore data scale effects. We apply DIAL on the final checkpoint, producing UserSim-160K-it1, UserSim-160K-it2, and UserSim-160K-it3 through successive refinement iterations. For comparison, we evaluate three zero-shot base models: GPT-4o (Hurst et al., 2024), Kimi-K2-Instruct (Bai et al., 2025), and Llama-3.3-70B-Instruct. Training hyperparameters are provided in Appendix C.2.
@@
For each configuration, we generate 2,000 simulated sessions: 1,000 unique session contexts drawn from real user profiles and conversation histories × 2 simulations per context, with early stopping when conversations naturally conclude. To isolate simulator quality effects, we use a fixed set of 14 chatbot models to simulate conversations.
@@ 4.3.1 Linguistic Features
Type-Token Ratio (TTR): The ratio of unique words (types) to total words (tokens) in a session, measuring lexical diversity. Higher TTR indicates greater vocabulary variety.
@@ 4.3.2 Behavioral Metrics
User Action Distribution: Conversations are annotated with task-relevant user actions using GPT-5-nano to classify each individual user message according to our action taxonomy (see Appendix D). Since individual messages can have multiple actions, the action distributions are not normalized.
Issue Rate Analysis: A critical requirement for realistic user simulation is the ability to expose chatbot failure modes. We evaluate simulators by measuring how well they replicate the rate and distribution of problematic conversational features (“issues”) that chatbots exhibit with real users. Issues are identified using GPT-5 with low reasoning effort. The analyzer identifies issues from a detailed taxonomy (35+ categories including invalidating responses, unclear wording, hallucinations, etc.) and highlights the specific messages where each issue occurred, enabling proper credit assignment (see Appendix E). We compare issue rates and distribution between simulated and real conversations to validate that our simulator accurately reflects the chatbot’s real-world performance and can reliably identify system weaknesses before deployment. We use KL divergence to measure distributional alignment:
@@
Similar to the discriminator, the issue analyzer examines whole sessions rather than isolated messages, which are only problematic when considered in context with the user’s situation and their subsequent replies. This enables numerous evaluation dimensions that would be impossible with isolated message analysis: detecting circular conversations, identifying premature topic changes, and assessing whether mental health progress occurs naturally over the conversation arc.
@@ 4.3.3 Discriminator Performance
We measure discriminator accuracy using fine-tuned discriminators trained on balanced datasets of real and simulated sessions from their corresponding simulators. Each discriminator is trained under identical conditions—1,600 training sessions and 400 validation sessions, with the same hyperparameters—ensuring that accuracy differences reflect simulator quality rather than training differences. Lower discriminator accuracy indicates a more realistic simulator.
@@ 5 Results
As a foundation for DIAL, we first establish that domain-specific SFT is necessary for realistic simulation. Fine-tuned simulators substantially outperform zero-shot models across all metrics. t-SNE visualization (Appendix B.1) shows that embeddings of conversations simulated by base models form distinct clusters separate from their real counterparts. The fine-tuned simulators’ outputs, on the other hand, form tight clusters defined by shared context, demonstrating effective grounding. Zero-shot models produce unrealistically compliant users, as evidenced by significantly lower issue rates (Figure 2) and overrepresentation of positive mental health indicators (Appendix B.3).
Table 1: Linguistic diversity features comparing real users with simulators. Values closer to real users indicate better linguistic realism. Best fine-tuned values (closest to real users) are in bold.
@@
As shown in Table 1, which compares linguistic features across simulators, some base models achieve diversity metrics closer to real users than fine-tuned models as SFT exhibits non-monotonic effects on lexical diversity. Bigram and TTR initially improve from 10K-40K samples as the model learns to use conversational context to improve diversity. However, mode collapse emerges from 40K-160K samples as continued training learns common patterns at the expense of rare variations. For behavioral alignment, fine-tuned simulators dramatically outperform zero-shot base models (see Appendix B.3). To evaluate how well simulators surface chatbot issues, we measure issue rates as the percentage of chatbot messages containing issues. Figure 2 shows that fine-tuned simulators produce rates of 3.07-3.31%, closely matching real conversations (3.02%), while zero-shot models show substantially lower rates of 0.98-2.54%. Detailed issue category discrepancies are provided in Appendix B.2.
@@
5.2 DIAL Training Improvements
Three iterations of DIAL produced progressive improvements:
@@ 5.2 DIAL Training Improvements
Linguistic Diversity: DIAL successfully addresses mode collapse by recovering lexical diversity while maintaining behavioral alignment, as models progressively approach real user levels across all diversity metrics. However, non-monotonic improvement across iterations indicates the need for careful hyperparameter tuning. We employ conservative dataset filtering to mitigate over-correction and maintain stability.
Distributional Alignment: DIAL produces modest improvements in action distribution alignment (see Appendix B.3). The limited magnitude of improvements suggests the base fine-tuned model already achieved strong alignment.
Discriminator Fooling: DIAL dramatically reduced discriminator accuracy across conversation turns (Table 2). All discriminators were trained under identical conditions with on-policy optimization. The dramatic performance degradation confirms that DIAL iterations remove systematic artifacts and produce strong distributional alignment throughout sessions.
Predictive Validity: To evaluate how well simulators enable reliable offline evaluation, we measure correlation between simulated and real issue rates across diverse chatbot models, and KL divergence between simulated and real issue category distributions. Table 3 shows that DIAL produces substantial improvements in both metrics. UserSim-160K-it2 achieves the strongest correlation ($r=0.818$), while UserSim-160K-it3 achieves the lowest distributional divergence (KL = 0.148). This provides compelling evidence that DIAL-trained simulators are significantly more realistic and effective at identifying chatbot problems in offline testing compared to their base fine-tuned counterparts. Notably, aggregating predictions across all simulator models achieves both a strong correlation and low distributional divergence, suggesting that ensembling multiple simulators could provide even more robust evaluation.
@@
A/B Test Replication: Beyond correlation, we validate predictive utility by replicating a recent model A/B test. All three models exhibited consistent directional trends between simulation and real data, and no model that failed in simulation improved in the wild (Appendix B.5).
@@ 6 Discussion
As context for the substantive contributions, note that zero-shot models show severe distributional biases, producing unrealistically compliant users who underrepresent negative and resistant behaviors. This finding confirms that domain-specific fine-tuning is a necessary foundation for realistic simulation, consistent with the broader literature on domain adaptation.
While SFT achieves strong behavioral alignment, it suppresses lexical diversity through mode collapse. DIAL, which avoids the training instability of prior policy gradient approaches (Li et al., 2017; Zhao et al., 2021) by employing a discrete iteration structure and preference-based optimization, recovers this diversity while maintaining behavioral fidelity.
DIAL produces simulators that can deceive increasingly sophisticated discriminators, demonstrating that realism requires both distributional accuracy and surface-level variation. Unlike traditional metrics that measure task success or correlation, discriminator accuracy directly quantifies distributional realism by testing whether conversations are distinguishable from real interactions.
@@ 7 Conclusion
This work presents DIAL (Direct Iterative Adversarial Learning), a framework for creating realistic user simulators that can reliably expose failure modes in multi-turn dialogue systems. Applied to mental health support, DIAL shows that domain-specific fine-tuning combined with iterative adversarial refinement produces simulators that closely replicate real user behavior across linguistic, semantic, and behavioral dimensions.
The key contributions of our work are: (1) the DIAL methodology, which uses discriminator-derived preference pairs with DPO instead of policy gradient methods, avoiding the training instability of prior adversarial dialogue approaches while recovering lexical diversity suppressed by supervised learning and improving distributional alignment; (2) demonstration that DIAL enables accurate identification of diverse failure modes in the mental health support domain; and (3) comprehensive validation showing that DIAL-trained simulators reliably predict real model performance, achieving strong correlation with real-world results and replicating A/B test trends.
These results show that DIAL is a promising methodology for creating realistic user simulators in multi-turn dialogue systems. By enabling rapid iteration and systematic exploration of system failure modes before deployment, such simulators can reduce development time, protect users from suboptimal experiences, and offer safe environments for RL-based optimization of conversation quality.
@@
Mode Collapse: Although DIAL substantially improves lexical diversity, fine-tuned simulators still exhibit lower phrase-level variety than real users. Techniques such as controlled generation or mixture-of-experts architectures may further enhance diversity preservation.
Domain Transfer: The evaluation in this study focuses on mental health support, a complex and open-ended TOD domain. Validating the methodology on traditional TOD benchmarks such as MultiWOZ (Budzianowski et al., 2018) and SGD (Rastogi et al., 2020), as well as in other complex domains such as tutoring and technical support, would establish broader applicability and help identify domain-specific challenges.
End-to-End TOD Training and Evaluation: Although the current simulators enable valid offline evaluation of chatbot failure modes, their application for training and evaluating end-to-end dialogue policies remains an area for future investigation. It is hypothesized that increasing simulator realism may not necessarily improve raw task success rates when evaluated by the simulator itself. Instead, success rates are expected to become more calibrated with real-user outcomes, as less realistic simulators often exhibit excessive compliance, leading to inflated success metrics. For instance, Cheng et al. (2022) train a user simulator with RL using task success as the reward, which encourages the simulator to become more compliant, contrary to the goal of surfacing system failures. Rigorous validation of this hypothesis requires access to real users for ground-truth comparison, a resource that is challenging to obtain at scale for most TOD benchmarks.
Human Evaluation: The evaluation of simulator realism in this study relies on automated metrics (discriminator accuracy, distributional alignment, predictive validity) rather than direct human judgment. Complementing these metrics with human evaluation of simulated conversation realism would provide additional validation and help identify failure modes that automated metrics may overlook.
@@
All data used in this research were collected from real user conversations with the mental health chatbot, with explicit user consent. Personally identifiable information was removed through automated script-based detection and removal of names, phone numbers, email addresses, and other identifiers in accordance with data protection policies and ethical guidelines for human subjects research. However, due to the sensitive nature of mental health data, the datasets used in this research cannot be shared, despite the intention to promote reproducibility and open science.
Realistic user simulators could potentially be misused to train systems optimized for engagement rather than user wellbeing, or to manipulate vulnerable populations. The simulators developed in this work are intended for system evaluation and development, enabling rapid iteration without exposing users to suboptimal experiences. It should be noted that simulators trained on real conversations may perpetuate biases present in the original dataset, potentially disadvantaging certain demographic groups. Simulators should therefore be deployed with appropriate oversight and ethical review. Additionally, while simulators can reduce the need for human testing in many scenarios, they should complement rather than entirely replace human feedback for validating novel interventions.
@@ B.3 User Action Distributions
For positive indicators (right), zero-shot models exhibit the opposite bias—substantially overrepresenting “Change Talk” ($\sim$40% vs. $\sim$15% in real data) and “Agrees with Chatbot.” This creates an artificially positive mental health support environment that fails to challenge the chatbot or surface issues. Fine-tuned simulators (UserSim variants) achieve close alignment with real sessions across both indicator types, with DIAL iterations showing incremental improvements in matching real distributions.
@@ B.4 Correlation with Production
DIAL substantially improves correlation with real issue rates, as shown in Table 3. UserSim-160K-it2 achieves $r=0.818$ compared to UserSim-160K-base’s $r=0.528$, demonstrating the value of DIAL refinement. However, we note that several models achieve strong correlations even without DIAL. This suggests that for rolled-out evaluation or RL training applications, ensembling multiple user simulator models could reduce noise and improve stability. Indeed, aggregating across UserSim-160K-base, it1, it2, and it3 achieves robust correlation ($r=0.801$) with low distributional divergence (KL = 0.152). The plot shown here highlights the best-performing model (UserSim-160K-it2) to demonstrate the upper bound of achievable predictive validity.
@@ C.2 Training Hyperparameters
Table 6 shows the hyperparameters used for SFT of the base user simulator and DPO during DIAL iterations.
@@ Appendix D User Action Categories
This appendix provides example user action categories from our taxonomy used to annotate and analyze user behavior in mental health support conversations. Actions are classified as positive mental health indicators (Good), negative indicators (Bad), or user features that should change over time. The full taxonomy contains 20+ categories; we highlight key examples below.
@@ Appendix E Issue Categories
This appendix provides example issue categories from our taxonomy used to evaluate chatbot quality. The full taxonomy contains 35+ categories organized into six main categories (User Distress, User Confusion, User Frustration, User Losing Trust, Chatbot Violates Policies, App Bugs); we highlight key examples below.